# Fundamental physical constants, operation of physical phenomena and entropy increase

### v5

@@ 5 Discussion and summary
The list of effects and phenomena in condensed matter physics discussed here can be probably extended significantly. The reason it is important to study how FPCs affect the operation and observability of different effects and phenomena in condensed matter is that this field is developed at the great level of detail. Physical models of condensed matter have been tested and cross-checked in numerous different experiments and simulations. In this sense, the insights into FPCs from condensed matter physics face less uncertainty and challenges as compared to models in high-energy physics and astrophysics (see Ref. [13] for a review of these challenges).
@@
Last but not least, deeper understanding the role of FPCs in condensed matter physics is important on its own. A fairly recent realisation that FPCs put lower or upper bounds on key condensed matter properties has deepened our understanding of how these properties operate and benefited our physical models. These properties include viscosity [17], diffusion and spin dynamics [18, 19, 20, 21, 22], electron and heat transport [23, 24, 25, 26, 27], thermal conductivity in insulators [28] and conductors [29], the speed of sound [30], elastic moduli [29, 31], flow in cells [32, 33], vibration frequency in solids and superconducting temperature [34]. Understanding how FPCs govern the observability and operation of entirely new effects and phenomena is another useful insight enhancing our understanding of condensed matter physics.

### v6

@@ V Discussion and summary
The list of effects and phenomena in condensed matter physics discussed here can be probably extended significantly. The reason it is important to study how FPCs affect the operation and observability of different effects and phenomena in condensed matter is that this field is developed at the great level of detail. Physical models of condensed matter have been tested and cross-checked in numerous different experiments and simulations. In this sense, the insights into FPCs from condensed matter physics face less uncertainty and challenges as compared to models in high-energy physics and astrophysics (see Ref. Adams (2019) for a review of these challenges).
@@
Last but not least, deeper understanding the role of FPCs in condensed matter physics is important on its own. A realization that FPCs put lower or upper bounds on key condensed matter properties has deepened our understanding of how these properties operate and benefited our physical models. These properties include viscosity Trachenko and Brazhkin (2020), diffusion and spin dynamics Luciuk et al. (2017); Sommer et al. (2011); Bardon et al. (2014); Trotzky et al. (2015); Enss and Thywissen (2018), electron and heat transport Hartnoll (2015); Mousatov and Hartnoll (2020); Behnia and Kapitulnik (2019); Nussinov and Chakrabarty (2022); Grozdanov (2021), thermal conductivity in insulators Trachenko et al. (2021) and conductors Brazhkin (2023), the speed of sound Trachenko et al. (2020), elastic moduli Brazhkin (2023); Trachenko (2023a), flow in cells Trachenko (2023b, c), vibration frequency in solids and superconducting temperature Trachenko et al. . Understanding how FPCs govern the observability and operation of entirely new effects and phenomena is another useful insight enhancing our understanding of condensed matter physics.